OpenStudy (anonymous):

help factor please x^2+2x+1 please show all work thanks!

OpenStudy (anonymous):

(x+1) (x+1) or (x+1)^2 If you want to check your answer, FOIL.

OpenStudy (anonymous):

ok thanks but how did you get it?

OpenStudy (anonymous):

\((a+b)^2=a^2+2ab+b^2\) with \(a=x, b=1\)

OpenStudy (anonymous):

(x+1) (x+1) you need a number that multiplied by itself gives you 1, and when added to each other gives you 2.

OpenStudy (anonymous):

oh wait sorry i forgot a negative sign in front of the x^2 so the equation is -x^2 + 2x + 1 but would that work for an equation with an axis of symmetry of 1

OpenStudy (anonymous):

1. Write two parenthesis ( ) ( ) 2. Since you have an x^2 that means you can easily put one x in each parentheses (x ) (x ) 3. Then find a two numbers that can give you the product of the last number and the same numbers have to give you the sum of the second number.....so since you have 2 and 1.. 1x1=1 and 1+1 = 2 therefore 4. (x+1) (x+1) and to verify that you use the foil method

OpenStudy (anonymous):

sorry but its -x^2

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Try the quadratic equation to find roots r1, r2. Then the factoring must be (x-r1)(x-r2). It is possible the roots are imaginary. That's fine, you can still write (x-r1)(x-r2).